# Rabi oscillations and entanglement between two atoms interacting by the Rydberg blockade and with a quantized radiation field studied by the Jaynes-Cummings Model

### v1

@@ 3 Number states of the field
|$\ket{\psi^{(n)}(t)}=\exp\left(-\frac{i}{\hbar}\hat{H}t\right)\ket{\psi^{(n)}(0)}$ (36)
$=\mu^{(n)}(t)\ket{\psi_{1}^{(n)}}+\nu^{(n)}(t)\ket{\psi_{2}^{(n)}}+\xi^{(n)}(t)\ket{\psi_{3}^{(n)}}$
@@
with |$\ket{\text{asym},n}=(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}})/\sqrt{2}$ defined in (17). This initial state corresponds to a maximally entangled Bell state of the two atoms and is an eigenstate of $\hat{H}$ in (9) with eigenvalue $E_{\text{asym}}^{(n)}=\hbar\omega_{f}n$ in (13). There is no transition of the field for this state. Thus,
@@
For the cases B and C discussed below, we quote the results for the coefficients $\mu^{(n)}(t)$, $\nu^{(n)}(t)$, $\xi^{(n)}(t)$ in (36) and the corresponding probabilities, leaving details to Appendices A and B, respectively. If not stated otherwise, first the result for general detuning $\Delta$ is given, followed by the result for $\Delta=0$.
@@
In this case, the two atoms are initially in their ground states and are exposed to the field in the photon number state $n+1$ at $t=0$. The coefficients in (36) are given by (see Appendix A.1)
$\mu_{B}^{(n)}(t)$ $=\frac{1}{\sqrt{2}}(-i)\sin(2\varphi_{n})\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\sin(\Omega_{n}t)$ (41)
@@
$\xi_{B}^{(n)}(t)$ $=\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\left[\cos^{2}(\varphi_{n})\exp(i\Omega_{n}t)+\sin^{2}(\varphi_{n})\exp(-i\Omega_{n}t)\right]$ (44)
@@
The probability to find the system in the complementary state |$\ket{\psi_{3}^{(n)}}$ is given by (see Appendix A.1) $P_{B}^{(n)}\big{(}\ket{\psi_{3}^{(n)}},t\big{)}=1-P_{B}^{(n)}(\text{sym},t)$.
@@
In view of Rydberg systems as potential building blocks of atom-light quantum interfaces in quantum networks [36, 39, 41], an important quantity is the entanglement between the two-atom system and the quantized field. Assuming zero detuning for simplicity ($\Delta=0)$, (46) and (42), (45) imply that at times $t$ for which $\lambda\sqrt{2(n+1)}\,t=\pi/4$ the state of the combined system (atoms + field) is given by
@@
with |$\ket{\text{sym},n}$ in (34) and $\psi_{3}^{(n)}=\ket{g,g,n+1}$ in (11). The state (49) corresponds to a maximally entangled state between the qubit encoded by the two-atom system ($\alpha$) as {|$\left\{\ket{g,g}_{\alpha},\ket{\text{sym}}_{\alpha}\right\}$ and the qubit encoded by the number states of the field ($f$) as {|$\left\{\ket{n}_{f},\ket{n+1)}_{f}\right\}$. If the photons of the field, after interacting with the two-atom system ($\alpha$) resulting in (49), are brought in contact with a second, separate two-atom system ($\beta$) initially prepared in the state |$\ket{g,g}_{\beta}$, then the second two-atom system will evolve to a superposition of |$\ket{g,g}_{\beta}$ and |$\ket{\text{sym}}_{\beta}$ which depends on the photon qubit {|$\left\{\ket{n}_{f},\ket{n+1)}_{f}\right\}$. That is, the second two-atom system ($\beta$) will be entangled with the first ($\alpha$) via the photons of the field. Although this scheme may be challenging to implement experimentally since it requires precise control of the photon number $n$, it illustrates how systems described by the extended Jaynes-Cummings model (4) can in principle be used to entangle stationary atomic systems via photons. In Section 4 we discuss the entanglement between a two-atom system and a quantum single-mode coherent field, which does not require precise control of individual photon number states |$\ket{n}$.
@@
In this case, one of the atoms is initially excited while the other is in the ground state, and the two atoms are exposed to the field in the photon number state |$\ket{n}$ at $t=0$. Since the two atoms are identical, without restriction we consider the case that atom $a$ is excited and atom $b$ is in the ground state at $t=0$. The coefficients in (36) are given by (see Appendix B)
@@
where the (+)(+)( + ) sign in (50), (51) belongs to $\mu_{C}^{(n)}(t)$ and the ()(-)( - ) sign belongs to $\nu_{C}^{(n)}(t)$. The angle $\varphi_{n}(\Delta)$ is given in (23). For the probabilities of |$\ket{\psi_{1}^{(n)}}$ and |$\ket{\psi_{2}^{(n)}}$ (given by $|\mu_{C}^{(n)}(t)|^{2}$ and $|\nu_{C}^{(n)}(t)|^{2}$, respectively) we only quote explicit formulae for zero detuning ($\Delta=0$) for simplicity; results for $\Delta\neq 0$ are shown in Figure 6 below.
@@ 4 Coherent states of the field
|$\ket{\alpha}=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\sum\limits_{m=0}^{\infty}\frac{\alpha^{m}}{\sqrt{m!}}\ket{m}\,,\,\,t=0\,,$ (55)
@@
|$\ket{\psi^{(\alpha)}(0)}$ =|$=\ket{g,g,\alpha}:=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\sum\limits_{m=0}^{\infty}\frac{\alpha^{m}}{\sqrt{m!}}\ket{g,g,m}$ (56)
$=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\left[\ket{g,g,0}+\sum\limits_{n=0}^{\infty}\frac{\alpha^{n+1}}{\sqrt{(n+1)!}}\ket{g,g,n+1}\right]\,,$
@@
|$\ket{\psi^{(\alpha)}(t)}=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\left[\exp(i\omega_{0}t)\ket{g,g,0}+\sum\limits_{n=0}^{\infty}\frac{\alpha^{n+1}}{\sqrt{(n+1)!}}\ket{\psi_{B}^{(n)}(t)}\right]\,,\,\,t>0\,,$ (57)
@@
$\hat{\rho}_{\text{atoms}}(t)$ $=\text{tr}_{f}\left\{\hat{\rho}(t)\right\}$ (61)
$=\sum\limits_{k=0}^{\infty}\langle k\,|\,\psi^{(\alpha)}(t)\rangle\langle\psi^{(\alpha)}(t)\,|\,k\rangle\,,$
@@
$P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$
$\quad=\sum\limits_{k=0}^{\infty}\langle\text{sym},k\,|\,\psi^{(\alpha)}(t)\rangle\langle\psi^{(\alpha)}(t)\,|\,\text{sym},k\rangle$
@@
$P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)=\langle g,g\,|\,\hat{\rho}_{\text{atoms}}(t)\,|\,g,g\rangle$
$\quad=\exp(-\bar{n})\left(1+\sum\limits_{m=1}^{\infty}\frac{\bar{n}^{m}}{m!}\left|\xi_{B}^{(m-1)}(t)\right|^{2}\right)$ (64)
@@
Figure 7 shows $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ in (59) for zero detuning ($\Delta=0$) and $\bar{n}=10$, $20$, and $50$ as a function of $t$ (blue lines). The oscillations of the probability undergo collapses and revivals on a time scale $\sim\lambda\sqrt{\bar{n}}$, similarly as for a single two-level atom interacting with a quantum single-mode coherent field [31, 13]. However, for the present two-atom system, the oscillations occur between |$\ket{g,g}$ and the maximally entangled Bell state |$\ket{\text{sym}}$ in (58); thus, the interaction of the two-atom system with a quantum single-mode coherent field entangles the two atoms, where the time-dependence of the states and the entanglement can be tuned by the average photon number $\bar{n}$ of the coherent field.
Figure 7: Time-dependence, in units of $\lambda^{-1}$, of the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ in (59) (blue lines) and the entanglement $\varepsilon(t)$ in (66) between the two-atom system and the quantum single-mode coherent field (orange lines) for $\omega_{f}=\omega_{0}=\lambda$ and (A) $\bar{n}=10$, (B) $\bar{n}=20$, and (C) $\bar{n}=50$. Part (C) highlights intervals in which $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ first collapses (I) and then remains constant at $1/2$ (II) before the first revival.
In view of the potential use of Rydberg systems for quantum state-transfer between atoms and photons in quantum information processing, it is of interest to study the entanglement of the two-atom system and the quantum coherent field. A measure of this entanglement is the quantity
$\varepsilon(t)=1-\text{tr}\{\hat{\rho}_{\text{atoms}}^{2}(t)\}\,,$ (66)
which varies between 0 when the atomic and field states are completely disentangled (where $\hat{\rho}_{\text{atoms}}$ describes a pure state) and $1/2$ when the atomic and field states are maximally entangled (where both eigenvalues of $\hat{\rho}_{\text{atoms}}$ are equal to $1/2$). In the basis {|$\left\{\ket{\text{sym}},\ket{g,g}\right\}$ $\hat{\rho}_{\text{atoms}}$ has the form
$\rho_{\text{atoms}}(t)=\begin{pmatrix}P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)&\gamma(t)\\[4.0pt] \gamma^{*}(t)&P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)\end{pmatrix}$ (67)
with
@@
This yields
$\text{tr}\{\hat{\rho}_{\text{atoms}}^{2}(t)\}=$ (69)
@@
$P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)^{2}+P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)^{2}+2\left|\gamma(t)\right|^{2}$
(explicit expressions for $\gamma(t)$ and $\left|\gamma(t)\right|^{2}$ are given in Appendix C).

### v2

@@ 3 Number states of the field
|$\ket{\psi^{(n)}(t)}=\exp\left(-\frac{i}{\hbar}\hat{H}t\right)\ket{\psi^{(n)}(0)}=\mu^{(n)}(t)\ket{\psi_{1}^{(n)}}+\nu^{(n)}(t)\ket{\psi_{2}^{(n)}}+\xi^{(n)}(t)\ket{\psi_{3}^{(n)}}$ (37)
@@
with |$\ket{\text{asym},n}=(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}})/\sqrt{2}$ defined in (17). This initial state corresponds to a maximally entangled Bell state of the two atoms and is an eigenstate of $\hat{H}$ in (9) with eigenvalue $E_{\text{asym}}^{(n)}=\hbar\omega_{f}n$ in (13). This atomic state is a dark state that does not couple to the field. Thus,
@@
For the cases B and C discussed below, we quote the results for the coefficients $\mu^{(n)}(t)$, $\nu^{(n)}(t)$, $\xi^{(n)}(t)$ in (37) and the corresponding probabilities, leaving details to A and B, respectively. If not stated otherwise, first the result for general detuning $\Delta$ is given, followed by the result for $\Delta=0$.
@@
In this case, the two atoms are initially in their ground states and are exposed to the field in the photon number state $n+1$ at $t=0$. The coefficients in (37) are given by (see A.1)
[…]
@@
$\xi_{B}^{(n)}(t)$ $=\exp\Big{[}-i\hskip 1.0pt\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}\hskip 1.0ptt\Big{]}\left[\cos^{2}(\varphi_{n})\exp(i\Omega_{n}t)+\sin^{2}(\varphi_{n})\exp(-i\Omega_{n}t)\right]$ (45)
@@
The probability to find the system in the complementary state |$\ket{\psi_{3}^{(n)}}$ is given by (see A.1) $P_{B}^{(n)}\big{(}\ket{\psi_{3}^{(n)}},t\big{)}=1-P_{B}^{(n)}(\text{sym},t)$.
@@
In view of Rydberg systems as potential building blocks of atom-light quantum interfaces, an important quantity is the entanglement between the two-atom system and the quantized field [37, 40, 42]. Assuming zero detuning for simplicity ($\Delta=0)$, (47) and (43), (46) imply that at times $t$ for which $\lambda\sqrt{2(n+1)}\,t=\pi/4$ the state of the combined system (atoms + field) is given by
@@
with |$\ket{\text{sym},n}$ in (35) and $\psi_{3}^{(n)}=\ket{g,g,n+1}$ in (11). The state (50) corresponds to a maximally entangled state between the qubit encoded by the two-atom system ($\alpha$) as {|$\left\{\ket{g,g}_{\alpha},\ket{\text{sym}}_{\alpha}\right\}$ and the qubit encoded by the number states of the field ($f$) as {|$\left\{\ket{n}_{f},\ket{n+1)}_{f}\right\}$. If the photons of the field, after interacting with the two-atom system ($\alpha$) resulting in (50), are brought in contact with a second, separate two-atom system ($\beta$) initially prepared in the state |$\ket{g,g}_{\beta}$, then the second two-atom system will evolve to a superposition of |$\ket{g,g}_{\beta}$ and |$\ket{\text{sym}}_{\beta}$ which depends on the photon qubit {|$\left\{\ket{n}_{f},\ket{n+1)}_{f}\right\}$. That is, the second two-atom system ($\beta$) will be entangled with the first ($\alpha$) via the photons of the field. Although this scheme may be challenging to implement experimentally since it requires precise control of the photon number $n$, it illustrates how systems described by the extended Jaynes-Cummings model (4) can in principle be used to entangle stationary atomic systems via photons. In Section 5 we discuss different types of entanglements between the atoms and a quantum single-mode coherent field, which does not require precise control of individual photon number states |$\ket{n}$.
@@
In this case, one of the atoms is initially excited while the other is in the ground state, and the two atoms are exposed to the field in the photon number state |$\ket{n}$ at $t=0$. Because the Hamiltonian (4) is symmetric under atom exchange, without restriction we assume that atom $a$ is excited and atom $b$ is in the ground state. The coefficients in (37) are given by (see B)
@@
where the (+)(+)( + ) sign in (51), (52) belongs to $\mu_{C}^{(n)}(t)$ and the ()(-)( - ) sign belongs to $\nu_{C}^{(n)}(t)$. The angle $\varphi_{n}(\Delta)$ is given in (23). For the probabilities of |$\ket{\psi_{1}^{(n)}}$ and |$\ket{\psi_{2}^{(n)}}$ (given by $|\hskip 1.0pt\mu_{C}^{(n)}(t)|^{2}$ and $|\nu_{C}^{(n)}(t)|^{2}$, respectively) we only quote explicit formulae for zero detuning ($\Delta=0$) for simplicity; results for $\Delta\neq 0$ are shown in Figure 6 below.
@@ 4 Coherent states of the field
|$\ket{\alpha}=\exp\left(-\frac{\left|\alpha\right|^{2}}{2}\right)\sum\limits_{m=0}^{\infty}\frac{\alpha^{m}}{\sqrt{m!}}\ket{m}\,,\,\,t=0\,,$ (56)
@@
|$\ket{\psi^{(\alpha)}(0)}$ =|$=\ket{g,g,\alpha}:=\exp\left(-\frac{\left|\alpha\right|^{2}}{2}\right)\sum\limits_{m=0}^{\infty}\frac{\alpha^{m}}{\sqrt{m!}}\ket{g,g,m}$ (57)
$=\exp\left(-\frac{\left|\alpha\right|^{2}}{2}\right)\left[\ket{g,g,0}+\sum\limits_{n=0}^{\infty}\frac{\alpha^{n+1}}{\sqrt{(n+1)!}}\ket{g,g,n+1}\right]\,,$ (58)
@@
|$\ket{\psi^{(\alpha)}(t)}=\exp\left(-\frac{\left|\alpha\right|^{2}}{2}\right)\left[\exp(i\omega_{0}t)\ket{g,g,0}+\sum\limits_{n=0}^{\infty}\frac{\alpha^{n+1}}{\sqrt{(n+1)!}}\ket{\psi_{B}^{(n)}(t)}\right]\,,\,\,t>0\,,$ (59)
@@
$\hat{\rho}_{\text{atoms}}(t)=\text{tr}_{f}\left\{\hskip 1.0pt\hat{\rho}(t)\right\}=\sum\limits_{k=0}^{\infty}\langle k\,|\,\psi^{(\alpha)}(t)\rangle\langle\psi^{(\alpha)}(t)\,|\,k\rangle\,,$ (63)
@@
$P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)=\sum\limits_{k=0}^{\infty}\langle\text{sym},k\,|\,\psi^{(\alpha)}(t)\rangle\langle\psi^{(\alpha)}(t)\,|\,\text{sym},k\rangle$ (64)
@@
$P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)=\langle g,g\,|\,\hat{\rho}_{\text{atoms}}(t)\,|\,g,g\rangle=\exp(-\bar{n})\left(1+\sum\limits_{m=1}^{\infty}\frac{\bar{n}^{m}}{m!}\left|\xi_{B}^{(m-1)}(t)\right|^{2}\right)$ (67)
@@
Figure 8 shows $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ in (61) and $P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)$ in (67) for zero detuning ($\Delta=0$) and $\bar{n}=10$, $20$, and $50$ as functions of time $t$ (orange and gray lines in parts (A), (C), (E), respectively). The oscillations of the probabilities undergo collapses and revivals on a time scale $\sim\lambda\sqrt{\bar{n}}$, similarly as for a single two-level atom interacting with a quantum single-mode coherent field [32, 13]. However, for the present two-atom system, the oscillations occur between |$\ket{g,g}$ and the maximally entangled Bell state |$\ket{\text{sym}}$ in (60); thus, the interaction of the two-atom system with a quantum single-mode coherent field entangles the two atoms, where the time-dependence of the states and the entanglement can be tuned by the average photon number $\bar{n}$ of the coherent field. The phenomenon of entanglement will be discussed in more detail in the following section.
5 Entanglement
Figure 7: Entanglements between various components of the total, tripartite system consisting of the two Rydberg-interacting atoms $a$, $b$ (purple) and the quantized radiation field $f$ (red). The total system is assumed to be in the pure state |$\ket{\psi^{(\alpha)}(t)}$ in (59). (A) Entanglement $E(ab-f)$ (blue) between the composite two-atom system $ab$ (enclosed by the ellipse) and the field $f$, and entanglement of formation $E(a-b)$ (green) between the two atoms $a$, $b$. (B) Entanglement $E(a-bf)$ (magenta) between atom $a$ and the composite (atom $b$)-field system $bf$ (enclosed by the ellipse).
In view of the potential use of Rydberg systems for quantum state transfer between atoms and photons in quantum information processing [20, 27, 5, 15, 33, 21, 29, 30, 31], it is of interest to study the entanglement between various components of the total system consisting of the two Rydberg-interacting atoms $a$, $b$ and the quantized radiation field (see Figure 7). For non-interacting atoms described by the Tavis-Cummings model, corresponding results were obtained in reference [28] (see Section 6).
To be specific, we assume that the field is in a quantum coherent state |$\ket{\alpha}$ given by (56). The time evolution of the entanglement between the atoms and the field depends crucially on the initial condition at $t=0$. For example, if the atomic system is initially in the dark state |$\ket{\text{asym}}=\frac{1}{\sqrt{2}}\left(\ket{r,g}-\ket{g,r}\right)$ (Section 3.1) the total system (atoms + field) will remain a product state |$\ket{\text{asym}}\otimes\ket{\alpha}$ for all times $t>0$ because the dark state does not couple to the field. This implies that the atoms will remain disentangled from the field. Conversely, if both atoms are initially in their ground state |$\ket{g}$ (Section 3.2) the atomic system will remain in the two-dimensional subspace perpendicular to |$\ket{\text{asym}}$ and spanned by |$\ket{g,g}$ and |$\ket{\text{sym}}=\tfrac{1}{\sqrt{2}}\left(\ket{r,g}+\ket{g,r}\right)$ for all times $t>0$ (compare Figure 3). In this case, the time evolution of the total system (atoms + field) is given by |$\ket{\psi^{(\alpha)}(t)}$ in (59) and discussed in Section 4. In what follows, we focus on this case because we expect that the entanglement between the atoms and the field will be strongest when the atomic system is perpendicular to the dark state |$\ket{\text{asym}}$. Results for other initial conditions, e.g., one atom excited and the other in the ground state, can be obtained using the results in Section 3.
5.1 Entanglement between the composite two-atom system $ab$ and the field $f$
@@
If the composite system is in a pure state |$\ket{\psi}$ with density operator $\hat{\rho}=\mathinner{|{\psi}\rangle\langle{\psi}|}$ a quantitative measure for the entanglement $E(A-B)$ between the two subsystems $A$ and $B$ is the von Neumann entropy of either of the two subsystems $A$ and $B$ [33, 73]:
$E(A-B)=S_{A}=-\text{tr}\left\{\hskip 1.0pt\hat{\rho}_{A}\log_{2}\hat{\rho}_{A}\right\}=S_{B}=-\text{tr}\left\{\hskip 1.0pt\hat{\rho}_{B}\log_{2}\hat{\rho}_{B}\right\}\,,$ (69)
[…]
Figure 8: Time-dependence, in units of $\lambda^{-1}$, of the atomic probabilities $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ and $P_{\text{atoms}}^{(\alpha)}\left(\ket{gg},t\right)$ in (61) and (67), respectively, and of the entanglements between various components of the total system consisting of the two atoms $a$, $b$ and the quantized radiation field $f$ (Section 5 and Figure 7). The total system is initially in the state |$\ket{g,g,\alpha}$ in (57). All plots are for $\omega_{f}=\omega_{0}=\lambda$, zero detuning ($\Delta=0$), and for the indicated values of $\bar{n}$. (A) Probabilities $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ (orange lines) and $P_{\text{atoms}}^{(\alpha)}\left(\ket{gg},t\right)$ (gray lines) of the atomic states |$\ket{\text{sym}}$ and |$\ket{gg}$, respectively. The sum of these probabilities is equal to 1 by normalization. (B) Entanglements for $\bar{n}=10$: Entanglement $E(ab-f;t)$ in (70) between the composite two-atom system $ab$ and the field $f$ (blue lines), entanglement $E(a-bf;t)$ in (73) between atom $a$ and the composite (atom $b$)-field system $bf$ (magenta lines), and entanglement of formation $E(a-b;t)$ in (78) between atoms $a$ and $b$ (green lines). The colors of the lines in (B) correspond to Figure 7. (C), (D) and (E), (F): same as (A), (B) for $\bar{n}=20$ and $\bar{n}=50$, respectively. Parts (E) and (F) highlight intervals in which $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ first collapses (I) and then remains constant at $1/2$ (II) before the first revival, showing the close correlation between the atomic probabilities and the entanglements.
We now apply these definitions to the total system consisting of the two atoms $a$, $b$ and the quantized radiation field $f$. We assume that the total system is in the pure state |$\ket{\psi^{(\alpha)}(t)}$ in (59) with density operator $\hat{\rho}(t)$ in (62) and consider the bipartition $ab\hskip 1.0pt|\hskip 1.0ptf$ of the total system into the composite two-atom system $ab$ (“atoms”) and the field $f$ (see Figure 7). Using (69) the entanglement $E(ab-f)$ between the two-atom system $ab$ and the field $f$ is given by
$E(ab-f)=S_{\text{atoms}}=-\text{tr}\left\{\hskip 1.0pt\hat{\rho}_{\text{atoms}}\log_{2}\hat{\rho}_{\text{atoms}}\right\}\,.$ (70)
The reduced density operator $\hat{\rho}_{\text{atoms}}=\text{tr}_{f}\left(\hskip 1.0pt\hat{\rho}\right)$ from (63) is represented in the basis {|$\left\{\ket{\text{sym}},\ket{gg}\right\}$ by the matrix
@@
$\rho_{\text{atoms}}(t)=\begin{pmatrix}P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)&\gamma(t)\\[4.0pt] \gamma^{*}(t)&1-P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)\hskip 1.0pt\end{pmatrix}\,\,,$ (71)
where $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ from (61) is the probability of finding the two atoms at $t>0$ in the state |$\ket{\text{sym}}$ in (60) and
[…]
An explicit expression for $\gamma(t)$ is given in C. $E(ab-f)$ in (70) varies between 0 when the atomic and field states are completely disentangled (where $\hat{\rho}_{\text{atoms}}$ describes a pure state) and 1 when the atomic and field states are maximally entangled (where both eigenvalues of $\hat{\rho}_{\text{atoms}}$ are equal to $1/2$). Figure 8 shows $E(ab-f)$ in (70) as a function of time $t$ for $\bar{n}=10$, $\bar{n}=20$, and $\bar{n}=50$ (blue lines in parts (B), (D), (F), respectively).
5.2 Entanglement $E(a-bf)$ between atom $a$ and the composite (atom $b$)-field system $bf$
Following the definitions in Section 5.1 we now consider the bipartition $a\hskip 1.0pt|\hskip 1.0ptbf$ of the total system into atom $a$ and the composite (atom $b$)-field system $bf$ (see Figure 7)  [^9] Switching atoms $a$ and $b$ does not change the results in this subsection because the Hamiltonian (4) is symmetric under atom exchange.. We again assume that the total system is in the pure state |$\ket{\psi^{(\alpha)}(t)}$ in (59). Using (69) the entanglement between the subsystems $a$ and $bf$ is given by
$E(a-bf)=S_{a}=-\text{tr}\left\{\hskip 1.0pt\hat{\rho}_{a}\log_{2}\hat{\rho}_{a}\right\}\,,$ (73)
where the reduced density operator $\hat{\rho}_{a}$ of atom $a$ can be obtained as the partial trace of $\hat{\rho}_{\text{atoms}}$ over atom $b$:
$\hat{\rho}_{a}=\text{tr}_{b,f}\left(\hskip 1.0pt\hat{\rho}\right)=\text{tr}_{b}\left(\hskip 1.0pt\hat{\rho}_{\text{atoms}}\right)\,.$ (74)
In the basis {|$\left\{\ket{g}_{a},\ket{r}_{a}\right\}$ the density operator $\hat{\rho}_{a}$ is given by the matrix
$\rho_{a}(t)=\begin{pmatrix}1-\frac{1}{2}P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)&\frac{1}{\sqrt{2}}\gamma^{*}(t)\\[4.0pt] \frac{1}{\sqrt{2}}\gamma(t)&\frac{1}{2}P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)\end{pmatrix}\,\,.$ (75)
Figure 8 shows $E(a-bf)$ in (73) as a function of time $t$ for $\bar{n}=10$, $\bar{n}=20$, and $\bar{n}=50$ (magenta lines in parts (B), (D), (F), respectively).
5.3 Entanglement of formation $E(a-b)$ between the two atoms $a$ and $b$
Consider again a composite quantum system $Q$ with Hilbert space $\mathbb{H}$ and a bipartition $A|B$ of this system into subsystems $A$ and $B$ as discussed at the beginning of Section 5.1. If the composite system is in a mixed state described by a density operator $\hat{\rho}$ rather than in a pure state |$\ket{\psi}$, the entanglement can no longer be obtained as the von Neumann entropy of the subsystems $A$ or $B$. Instead, a quantitative measure for the entanglement $E(\hskip 1.0pt\hat{\rho})$ is now given by the entanglement of formation defined as the minimum average entanglement entropy over all pure state decompositions of $\hat{\rho}$, i.e., minimized over all ensembles of states |$\ket{\psi_{i}}$ with probabilities $p_{i}$ such that $\hat{\rho}=\sum_{i}p_{i}\mathinner{|{\psi_{i}}\rangle\langle{\psi_{i}}|}$ [74, 75, 76, 73]. The entanglement of formation is efficiently computable for two qubits [76], in our case the two two-state atoms $a$ and $b$. More general multipartite entanglement measures for pure and mixed states are reviewed in [73].
[…]
$E(\hskip 1.0pt\hat{\rho})=h\left(\frac{1}{2}+\frac{1}{2}\sqrt{1-C(\hskip 1.0pt\rho)^{2}}\right)\,\,,$ (76)
where $h(x)=-x\log_{2}x-(1-x)\log_{2}(1-x)$.
We now apply these definitions to the mixed state of the composite two-atom system $ab$ (“atoms”) represented by the reduced density operator $\hat{\rho}_{\text{atoms}}(t)=\text{tr}_{f}\left\{\hskip 1.0pt\hat{\rho}(t)\right\}$ given in (63). This results in the entanglement of formation $E(a-b)$ of the two-atom system $ab$, corresponding to the entanglement of atom $a$ with atom $b$ (see Figure 7). In the standard two-qubit basis quoted above the density operator $\hat{\rho}_{\text{atoms}}$ is represented by the matrix
$\rho_{\text{atoms}}(t)=\begin{pmatrix}P_{gg}(t)&\frac{1}{\sqrt{2}}\gamma^{*}(t)&\frac{1}{\sqrt{2}}\gamma^{*}(t)&0\\[4.0pt] \frac{1}{\sqrt{2}}\gamma(t)&\frac{1}{2}P_{\text{sym}}(t)&\frac{1}{2}P_{\text{sym}}(t)&0\\[4.0pt] \frac{1}{\sqrt{2}}\gamma(t)&\frac{1}{2}P_{\text{sym}}(t)&\frac{1}{2}P_{\text{sym}}(t)&0\\[4.0pt] 0&0&0&0\end{pmatrix}\,\,.$ (77)
Here, we use the abbreviations $P_{gg}(t)=P_{\text{atoms}}^{(\alpha)}\left(\ket{gg},t\right)$ and $P_{\text{sym}}(t)=P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ for ease of notation, where $P_{gg}(t)+P_{\text{sym}}(t)=1$ by normalization. The rightmost column and the bottom row of the matrix (77) are filled with zeros because the doubly excited state |$\ket{rr}$ is not available to the system due to the Rydberg blockade between the atoms. We find that the matrix $\rho_{\text{atoms}}\cdot\tilde{\rho}_{\text{atoms}}$ has only 1 nonzero eigenvalue given by $P_{\text{sym}}^{\,2}$ resulting in $C(\hskip 1.0pt\rho_{\text{atoms}})=P_{\text{sym}}=P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}}\right)$. That is, the concurrence is precisely equal to the probability of finding the two atoms in the state |$\ket{\text{sym}}$ in (60). Using (76) we obtain the entanglement of formation as a function of time $t$:
$E(a-b;t)=h\left(\frac{1}{2}+\frac{1}{2}\sqrt{1-\left[P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)\right]^{2}}\right)\,\,.$ (78)
Clearly, $E(a-b;t)$ is entirely determined by the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ (Figure 9). The limiting behaviors are $E(a-b)\sim P_{\text{sym}}^{\,2}$ for $P_{\text{sym}}\to 0$ (with a logarithmic correction) and $E(a-b)\to 1$ linearly for $P_{\text{sym}}\to 1$.
Figure 9: Entanglement of formation $E(a-b)$ between the two atoms $a$ and $b$ in (78) as a function of the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}}\right)$ of finding the two atoms in the state |$\ket{\text{sym}}$ in (60).
Figure 8 shows $E(a-b;t)$ in (78) as a function of time $t$ for $\bar{n}=10$, $\bar{n}=20$, and $\bar{n}=50$ (green lines in parts (B), (D), (F), respectively).